OpenStudy (anonymous):

Jon and Lexi are playing a marble game. According to their rules, 6 points are awarded for each red marble collected and 2 points are awarded for each blue marble collected. Which expression represents Jon's total score if he has r red marbles and b blue marbles? 6r - 2b 6r + 2b 6r - 4b 8rb

OpenStudy (anonymous):

@kcurri wouldn't this be a

OpenStudy (anonymous):

I think it would be B. Because 6r + 2b is 6 = red and 2 = blue. and is says his total score. Do you understand? :)

OpenStudy (anonymous):

yes and everybody is telling me I got this one answer wrong can you check it?

OpenStudy (anonymous):

ok I can try!

OpenStudy (anonymous):

Step 1 3x-4x+1=4 Step 2 x+1=4 Step 3 x+1-1=4-1 Step 4 x=3

OpenStudy (anonymous):

so to check this you would wright it out. 3*3 - 4*3 + 1 = 4 9-12+1 = 4 3+1=4 so yes you are right! x=3! Good Job!

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Welcome! ;)